Science Report – Nov 24th

Crew 216 Science Report 24-Nov-2019
Sol 0
Submitted by: Crew Researcher Whittle

ATTN: More detail to follow on Sol 1 in conjunction with the finalization of the mission plan.

1) Psych Scale (TAMU IRB2019-1042D) – Pre-mission surveys collected from all crew on Sol -1, next collection on D+5. Blood Pressure collection from 1 subject at 0900T each morning, a first reading taken this morning.

2) Medical Casts – Continuation of Crew 215 research planned. 2 EVAs requested for Sol 1 to collect gypsum samples.

Operations Report – Nov 24th

Crew 216 Operations Report 24-Nov-19 Amended
SOL: 0
Name of person filing report: Mike Lawson

Non-nominal systems:
Notes on non-nominal systems: N/A

Generator: Run
Hours run: 12
From what time last night: 1909
To what time this morning: 0709
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 56

Diesel Reading – 44
Station Propane Reading – 62
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1461672
Water (static tank) (gal): 21
Static to Loft Pump used: Yes
Water in Green Hab (gal): 280
Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 113.6
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 56
Currently charging: Yes

Opportunity rover used: Yes
Hours: 68.4
Beginning charge: 100
Ending charge: 60
Currently charging: Yes

Curiosity rover used: Yes
Hours: 118.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 66
Currently charging: Yes

Notes on rovers: All OK

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where? Returned Crew 215 to Grand Junction.

General notes and comments:
Questions, concerns and requests to Mission Support:

Summary of Internet: Functional
Summary of suits and radios: Functional
Summary of Hab operations: Functional
Summary of GreenHab operations: Changed shade cloth.
Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Solar is not tracking on the sun.
Summary of health and safety issues: Nominal

Commander Report – November 24th

Crew 216 Commander Report 24-NOV-2019
Sol: 0
Summary Title: Turnover and training day
Author’s name: Marc Levesque

Today Crew 216 said good-bye to Crew 215 and completed its training on rover and EVA operations and a final briefing from the Mission Director.

The crew is settling in very well, with all hands pitching in on various tasks during the day and anticipating starting sim tomorrow at 0700.

Two photos to follow with the Sunrise on Sol 0 as our Photo of the day.

Longer Commander Reports will be forthcoming, as will more photos.

Greenhab Report – Nov 24th

Crew 216 GreenHab – 24-NOV-2019
GreenHab Officer: Michael Ho

Environmental control: Door left open 12.40 pm, door closed 4.50 pm.

Average temperature: 24C

Hours of supplemental light: Automatic system

Daily water usage for crops: 16L (2 blue cans)
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 12:40 pm; 4.50 pm

Changes to crops: 6 additional species planted as seed on 23/11, details listed below. 1 avocado seed planted today.

Summer shade cloth taken down, 50% shade cloth hung, green shelves rotated 90 degrees to a new position.

The following planted seeds have germinated as of today – oregano, chives, marjoram, lemon balm, basil, thyme, sage, coriander, carrots, radish, spinach (2 species), rocket, tomatoes (2 species), cucumber, onion, snow peas.

SEASONAL PLANTING HISTORY
Established plants introduced 03/11/2019

[3x] strawberry plants (Everbearing), in a medium metal planter
[1x] spicy orange thyme (thymus fragantissimus), small terra-cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra-cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
[2x] small Aloe Vera, in a medium metal planter, frost-damaged (removed 1311219)

Seeds Planted 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officinalis
Sage – Salvia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted 10/11/2019 (Large metal planting tub)
Coriander – Coriandrum Sativum

Seeds Planted in12/11/2019 (2 smaller planting tubs)
Carrots ‘Purple Sun F1’
Radish ‘icicle’

Seeds Planted 13/11/2019
Perpetual Spinach, Rocket, Spinach ‘Lakeside F1’in medium metal planter.
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.

Seeds Planted 14/11/2019
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.
Onion (Evergreen White Bunching)
Cucumbers (Patio Snackers F1) in a hanging pot
Pepper (Mini Bell Color Mix) in small pots for transplantation.

Seeds Planted 15/11/2019
Acorn Pumpkin in a small pot for future transplantation. The seed taken from a pumpkin brought to the Hab at the start of rotation as a vegetable.
Snow Peas in 2 hanging pots.

Bulbs planted 19/11/2019
25 white Narcissus bulbs (Paperwhites ‘Ziva’)
Planted into 3 small terra cotta pots & 2 long planter boxes.

Seed planted 21/11/2019
1 apple seed (green apple) from fruit brought to the campus early in the mission.

Seeds Planted 22/11/2019
Snow Peas in 3 hanging pots.

Seeds Planted 23/11/2019
Zucchini in 2 large black pots
Chili Pepper ‘Shishito’ in 2 2 gallon black pots
Sweet Greenfield Pepper ‘California Wonder’ in 2 2 gallon black pots
Swiss Chard ‘Flamingo Pink’ in 4 2 gallon black pots

Seed Planted 24/11/2019
1 avocado seed in a small black pot

Three self-seeded tomato plants in large blue tub transplanted to 2-gallon pots

OTHER NOTES:

Partly used packed of seed have been stored in the provided tin to protect from vermin.

Harvest: None

Support/supplies needed: It would be desirable for more wire hanging points to be installed on the roof frame for white IKEA hanging pots.

[end]

Sol Summary – Nov 24th

Crew 216 Sol Summary Report 24-NOV-2019
Sol: 0
Summary Title: Training Day
Author’s name: Marc Levesque, Commander

Mission Status: Active

Sol Activity Summary: Today was a training day in which the crew learned about rover operations, EVA equipment, and procedures, and overall operations of the Hab and sim requirements.

Anomalies in work: N/A.

Weather: Freezing overnight, light winds, clear skies.

Crew Physical Status: Healthy

EVA: N/A

Reports to be filed: Operations Report, GreenHab Report, Science Report, EVA Requests, HSO Pre-Mission Checklist, and Commander Report.

Support Requested: N/A